If you had 0.956 moles of a nonelectrolyte solute in 345 grams of water. What is the new freezing point? (The Kf of water is 1.86 °C/m ) *
1
Expert's answer
2020-06-02T14:06:33-0400


solution molarity

"M=\frac{n}{V}=\frac{0.956}{0.345}=2.77"


freezing point depression

"\Delta Tf=Kf\times M=1.86\times2.77=5.15"

freezing point

"\Delta Tfsol=0-5.15=-5.15" °C